# MoD2T:Model-Data-Driven Motion-Static Object Tracking Method

### v1

@@ Abstract
The domain of Multi-Object Tracking (MOT) is of paramount significance within the realm of video analysis. However, both traditional methodologies and deep learning-based approaches within this domain exhibit inherent limitations. Deep learning methods driven exclusively by data exhibit challenges in accurately discerning the motion states of objects, while traditional methods relying on comprehensive mathematical models may suffer from suboptimal tracking precision. To address these challenges, we introduce the Model-Data-Driven Motion-Static Object Tracking Method (MoD2T). We propose a novel architecture that adeptly amalgamates traditional mathematical modeling with deep learning-based MOT frameworks, thereby effectively mitigating the limitations associated with sole reliance on established methodologies or advanced deep learning techniques. MoD2T’s fusion of mathematical modeling and deep learning augments the precision of object motion determination, consequently enhancing tracking accuracy. Our empirical experiments robustly substantiate MoD2T’s efficacy across a diverse array of scenarios, including UAV aerial surveillance and street-level tracking. To assess MoD2T’s proficiency in discerning object motion states, we introduce MVF1 metric. This novel performance metric is designed to measure the accuracy of motion state classification, providing a comprehensive evaluation of MoD2T’s performance. Meticulous experiments substantiate the rationale behind MVF1’s formulation. To provide a comprehensive assessment of MoD2T’s performance, we meticulously annotate diverse datasets and subject MoD2T to rigorous testing. The achieved MVF1 scores, which measure the accuracy of motion state classification, are particularly noteworthy in scenarios marked by minimal or mild camera motion, with values of 0.774 on the KITTI dataset, 0.521 on MOT17, and 0.827 on UAVDT.
@@ I Introduction
MULTI-OBJECT tracking (MOT) stands as a pivotal cornerstone in the realm of video analysis, capturing the dedicated attention of scholars and researchers alike. The inception of MOT was bound by the limitations inherent in detection techniques, prompting initial pursuits that encompassed methodologies like image differencing, Kalman filtering, and the Hungarian algorithm, laying the foundational bedrock of tracking frameworks[1]. Progressing forward, the evolution witnessed the emergence of motion detection techniques, inclusive of the application of frame differencing. These methods, underpinned by statistical models exploiting pixel or feature-based statistical properties, provided the groundwork for models detecting anomalies or changes. Furthermore, statistical model-centric strategies materialized, exemplified by Gaussian mixture models [2] and adaptive background modeling [3]. It remains conspicuous that traditional non-deep learning methods, fixated solely on tracking moving objects, often yield suboptimal outcomes across diverse scenarios. This shortcoming, attributed to the intricate nature of image complexities, inevitably curtails the precision of these conventional methodologies. In the ever-evolving landscape of computer vision, a profound transformation has unfolded, primarily steered by the pervasive embrace of deep learning. This transformative wave finds its embodiment in two prominent methodologies: [4, 5, 6] and joint-detection-association[7, 8, 9, 10]. Tracking-by-detection intricately harnesses the power of deep learning models for precise object detection, seamlessly integrating them with sophisticated techniques like Kalman filtering to pave the way for seamless object tracking. On the flip side, joint-detection-association unfolds as an end-to-end integrated tracking network. The deep learning paradigm brings a tapestry of advantages to the forefront in the realm of MOT. It endows the system with unparalleled capabilities for feature learning and representation, allowing for a nuanced understanding of the semantic intricacies of tracked objects. Deep learning’s forte lies in its adeptness at modeling context effectively, enabling a holistic comprehension of the interplay between tracked objects and their surroundings[11]. The scalability of deep learning models in handling vast amounts of data adds another layer of strength, a crucial facet given the large-scale nature of MOT tasks. Furthermore, the principle of transfer learning accelerates the adaptation of models from one task to another, a vital asset in the ever-evolving MOT landscape. Adaptability emerges as a hallmark of deep learning, particularly in navigating the complexities of dynamic scenes marked by occlusions and intricate motion patterns. This innovative era has not only witnessed the ascendancy of tracking-by-detection but has also translated into remarkable successes across a diverse spectrum of applications[12, 13, 14]. However, these neural network-based methods heavily rely on object appearance features during implementation. Consequently, they often track all objects within an image, deviating from traditional methods that specifically target moving objects. Environments such as vehicle monitoring, battlefield operations, and autonomous driving typically require a focus on moving objects. In these scenarios, there is a need for information regarding the motion of objects relative to the ground, a challenge current methods struggle to address. Furthermore, as many scenarios utilize single RGB cameras rather than stereo or RGB-D cameras, discerning object motion becomes even more challenging. Addressing this issue, this paper introduces a straightforward yet practical approach for determining object motion, adaptable across various environments.
@@
Traditional non-deep learning methods adeptly track moving targets, harnessing the physical and mathematical facets of images and relying on comprehensive mathematical models. While they might exhibit reduced accuracy in select cases, their efficacy extends across almost any environment. In contrast, deep learning-driven MOT frameworks pivot on data-driven principles. Though their precision is notable, they may falter when confronted with limited or inadequate data, leading to inaccuracies in tracking object motion estimation. To transcend the limitations inherent in both deep learning algorithms and traditional methods, we present a tracking approach that melds conventional mathematical modeling with data-driven deep learning methodologies. In essence, this paper unfolds with three core contributions: 1) proposing metric, MVF1, for evaluating the model’s ability to determine the relative motion status of objects with respect to the ground. Experimental validation of MVF1 is conducted, providing evidence for its rationality; 2) introducing a plug-and-play method for discerning the motion attributes of tracked objects; and 3) proposing a model-data-driven tracking framework(MoD2T) that harmonizes traditional methods with MOT tracking frameworks, enabling the delineation of object motion states.
@@ II-A Motion Objects Detection Methods
Traditional MOT methods primarily focus on identifying moving targets, and the subsequent tracking process is similar to the SORT[4]. In the early stages, image change detection methods were used for this purpose. The development of image change detection methods can be traced back several decades. The earliest methods were based on pixel value comparison to detect changes [15, 1, 16]. These methods were simple and intuitive but sensitive to complex scenes and noise. Subsequently, methods based on pixel-level differences received extensive research attention. These methods use pixel-level statistics or thresholds to detect changes, such as pixel ratio [17] and statistical parameters [18]. While these methods are robust to noise, they still struggle with complex textures and lighting variations [19]. Before the surge of deep learning, various Gaussian background model were the state-of-the-art in this field. Specifically, for a background image, the grayscale values of specific pixel points in the temporal domain follow a Gaussian distribution. In other words, for a background image $B$, the grayscale value $x$ of each point $(u,v)$ satisfies $I(x)\sim N\left(\mu,\sigma^{2}\right)$ as shown in (1):
@@
If the condition is met, the point is considered a background point; otherwise, it is considered a foreground point. Here, $T$ is a constant threshold representing the grayscale value at a given point in image $IG$, and $IB$ is similar in nature. In the field of image processing, where deep learning methods dominate, motion detection methods primarily adopt sparse feature tracking[20, 21, 22] or dense optical flow [23, 24, 25]. The former is widely used in SLAM, while the latter is commonly applied for overall motion estimation in images. Although each method has been proven effective for its respective applications, neither can fully model the motion in videos: optical flow cannot capture long-term motion trajectories, and sparse tracking cannot model the motion of all pixels. Recently, someone proposed a method combining the Optical flow with some other modes [26], and achieved good results. However, these methods still have certain limitations. Sparse feature tracking involves random selection of feature points, which may result in tracking objects with too few or no features, or unnecessary computational overhead if too many features are selected. The limitations of optical flow methods based on CNN [27, 28] or generative adversarial networks (GAN)[29, 30] include current challenges regarding their accuracy and universality. Traditional optical flow methods, on the other hand, exhibit high tracking accuracy in current MOT tracking frameworks. Of course, in recent years, several models based on deep learning for detecting changes in images have been proposed[31, 32, 33]. These methods have been widely used in fields such as remote sensing and change detection. However, due to the fact that their working principle relies on static image comparison, these methods are no longer suitable for motion object detection.
II-B MOT Methods
The developmental trajectory of MOT methodologies has been expounded upon in the preceding text and will not be extensively reiterated here. Currently, the available MOT approaches capable of distinguishing object motion exhibit limited proliferation. A significant portion of these approaches do not solely rely on RGB image information, but rather necessitate supplementary sensing modalities such as RGB-D cameras[34] or various alternative sensors like LIDAR[20]. The conventional MOT approach employed within this paoer shall be elucidated in the appendix. As for the deep learning-driven MOT method, this paper intends to embrace the tracking-by-detection paradigm. For the object detection stage, YOLOv8[35] will be used. YOLOv8 is a fast and accurate object detection method that divides the image into grids and predicts bounding boxes and their corresponding classes for object detection. For the subsequent tracking stage, StrongSORT [36] will be employed. StrongSORT is a deep learning-based object tracking method that combines appearance features and motion information to achieve accurate multiple object tracking. The tracking framework, YOLOv8+StrongSORT, combines the YOLOv8 object detector with the StrongSORT tracker, resulting in powerful tracking performance. YOLOv8 can rapidly and accurately detect targets in video frames and provide bounding box locations and class information for each detection. StrongSORT utilizes these detection results and combines them with the appearance features and motion information of the targets. It performs data association and trajectory prediction for target tracking and re-identification. Additionally, the paper has also experimented with other tracking methods, such as YOLOv8+DeepOCSORT[6], to explore multiple tracking approaches. Our method can be integrated with different MOT methods, and this article only selected two.
@@ II-B MOT Methods
Figure 1: Flow chart of motion judgment. The background feature extraction process is shown in Fig 2. And, in this paper, The frame rate difference $n$ for calculating distance and background similarity is set to 3. The specific meaning of the formula and the specific meanings of each parameter can be found in section III. The main purpose of this process is to calculate two indicators, one is the distance of the object’s movement in the image, and the other is the similarity of the object’s background in the image. Finally, these two indicators are combined to determine whether the object is moving relative to the ground.
II-C Fusion Method
The fusion of tracking methods involves combining multiple individual tracking methods or models to improve tracking performance or address various challenges. This approach can leverage the strengths of different methods, mitigate the limitations of individual methods, and provide more robust and accurate tracking results. One common method of tracking method fusion is decision-level fusion [37, 38, 39]. This method determines the final tracking result by weighting or voting the outputs of multiple trackers. The weights of each tracker can be dynamically adjusted based on its tracking performance in previous frames, giving greater weight to trackers with better performance. Another common fusion method is feature-level fusion[40, 41]. In this approach, multiple trackers extract feature representations of the target and fuse these features. Fusion can be as simple as averaging or weighted averaging, or it can involve more complex model fusion methods such as convolutional neural networks or support vector machines. Additionally, there are information fusion-based methods[42, 43] that involve fusing motion information, contextual information, or external prior knowledge. These methods improve tracking performance by combining information from multiple sources. For example, the results of multiple trackers can be fused by incorporating the target’s motion model, scene context, or appearance model. Overall, the fusion of tracking methods offers a way to enhance tracking performance by integrating multiple methods or models, leveraging their strengths, and taking advantage of different sources of information.
III Motion Judgment
@@
III-A Moving Distance
MOT methods are employed to discern the motion trajectories of objects within video sequences. In scenarios involving camera motion, the application of the enhanced correlation coefficient maximization(ECC) [44] model enables robust camera motion correction, allowing us to reliably ascertain the relative inter-frame displacement of objects. This computed displacement serves as a foundational metric for characterizing object motion. We employ the Mahalanobis distance to calculate the motion distance $D$, as shown in (1):
@@ III-A Moving Distance
Here, $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$ represent the center coordinates of the two bounding boxes, and $(x_{1},y_{1},w_{1},h_{1})$ and $(x_{2},y_{2},w_{2},h_{2})$ denote the coordinates and dimensions of the two bounding boxes. The rationale behind selecting Mahalanobis distance lies in its ability to account for the covariance structure within the data. This becomes particularly crucial when dealing with multidimensional data, as it considers both the variances and covariances among the different dimensions. Notably, we conducted experiments with alternative distance metrics, and Mahalanobis distance demonstrated superior performance in capturing the underlying data structure, especially in handling the intricacies of motion patterns represented by the bounding boxes. This choice significantly contributes to the accuracy and reliability of our motion distance calculations, ensuring robustness in the face of varying data distributions and dimensions.
III-B Background Feature
We performed feature extraction on the background of the tracked object. The specific approach involved segmenting the surrounding background and then combining it. We utilized a pre-trained OSNet [45] to extract appearance features from this background. The procedure is illustrated in the Fig.2. Empirical findings have substantiated the enhancement in accuracy resulting from the incorporation of this procedural step. The discernible rise in MVF1 scores, as illustrated in Tables IV, V, and VI, under the condition $\lambda=best$, underscores its efficacy in contrast to cases where $\lambda$ is set to 0. The subsequent section will provide a comprehensive explication of the variable $\lambda$, elucidating its impact on MVF1 scores. The optimal value for $\lambda$, denoted as $best$, is determined through systematic parameter tuning, with specific values tailored for each dataset. For brevity, explicit enumeration of these dataset-specific $best$ values is omitted in this context.
III-C Motion Judgment

### v2

# Motion State: A New Benchmark Multiple Object Tracking

@@ Abstract
In the realm of video analysis, the field of multiple object tracking (MOT) assumes paramount importance, with the motion state of objects—whether static or dynamic relative to the ground—holding practical significance across diverse scenarios. However, the extant literature exhibits a notable dearth in the exploration of this aspect. Deep learning methodologies encounter challenges in accurately discerning object motion states, while conventional approaches reliant on comprehensive mathematical modeling may yield suboptimal tracking accuracy. To address these challenges, we introduce a Model-Data-Driven Motion State Judgment Object Tracking Method (MoD2T). This innovative architecture adeptly amalgamates traditional mathematical modeling with deep learning-based multi-object tracking frameworks. The integration of mathematical modeling and deep learning within MoD2T enhances the precision of object motion state determination, thereby elevating tracking accuracy. Our empirical investigations comprehensively validate the efficacy of MoD2T across varied scenarios, encompassing unmanned aerial vehicle surveillance and street-level tracking. Furthermore, to gauge the method’s adeptness in discerning object motion states, we introduce the Motion State Validation F1 (MVF1) metric. This novel performance metric aims to quantitatively assess the accuracy of motion state classification, furnishing a comprehensive evaluation of MoD2T’s performance. Elaborate experimental validations corroborate the rationality of MVF1. In order to holistically appraise MoD2T’s performance, we meticulously annotate several renowned datasets and subject MoD2T to stringent testing. Remarkably, under conditions characterized by minimal or moderate camera motion, the achieved MVF1 values are particularly noteworthy, with exemplars including 0.774 for the KITTI dataset, 0.521 for MOT17, and 0.827 for UAVDT.
@@ I Introduction
The determination of whether an object is in motion, i.e., its motion state, is deemed to hold significant practical implications across various domains. For instance, in the realm of autonomous driving, moving objects often warrant greater attention than stationary ones, as foreknowledge of the motion state of objects on the road surface could potentially aid in autonomous driving decision-making. Similarly, in the field of remote sensing, the focus on moving objects supersedes that on stationary ones. Given that discerning the motion state of objects necessitates knowledge of their continuous positional coordinates, we posit that this issue constitutes a subtask within the domain of multiple object tracking (MOT).
MOT stands as a pivotal cornerstone in the realm of video analysis, capturing the dedicated attention of scholars and researchers alike due to its fundamental importance in numerous real-world applications. MOT originated from the necessity to overcome the limitations inherent in traditional detection techniques, prompting initial pursuits that encompassed methodologies like image differencing, Kalman filtering, and the Hungarian algorithm, thereby laying the foundational bedrock of tracking frameworks [1]. Subsequent advancements witnessed the emergence of motion detection techniques, including the application of frame differencing, and statistical model-centric strategies, exemplified by Gaussian mixture models [2] and adaptive background modeling [3]. However, it remains conspicuous that traditional non-deep learning methods, fixated solely on tracking moving objects, often yield suboptimal outcomes across diverse scenarios due to the intricate nature of image complexities.
In the ever-evolving landscape of computer vision, a profound transformation has unfolded, primarily steered by the pervasive embrace of deep learning. This transformative wave finds its embodiment in two prominent methodologies: tracking-by-detection and joint-detection-association. Tracking-by-detection intricately harnesses the power of deep learning models for precise object detection, seamlessly integrating them with sophisticated techniques like Kalman filtering to pave the way for seamless object tracking [4, 5, 6]. On the flip side, joint-detection-association unfolds as an end-to-end integrated tracking network [7, 8, 9, 10, 11]. The deep learning paradigm brings a tapestry of advantages to the forefront in the realm of MOT, endowing the system with unparalleled capabilities for feature learning and representation, allowing for a nuanced understanding of the semantic intricacies of tracked objects [12, 13].
Deep learning’s forte lies in its adeptness at modeling context effectively, enabling a holistic comprehension of the interplay between tracked objects and their surroundings [12]. The scalability of deep learning models in handling vast amounts of data adds another layer of strength, a crucial facet given the large-scale nature of MOT tasks. Furthermore, the principle of transfer learning accelerates the adaptation of models from one task to another, a vital asset in the ever-evolving MOT landscape.
Adaptability emerges as a hallmark of deep learning, particularly in navigating the complexities of dynamic scenes marked by occlusions and intricate motion patterns. This innovative era has not only witnessed the ascendancy of tracking-by-detection but has also translated into remarkable successes across a diverse spectrum of applications [14, 15, 16].
However, these neural network-based methods heavily rely on object appearance features during implementation. Consequently, they often track all objects within an image, deviating from traditional methods that specifically target moving objects. Environments such as vehicle monitoring, battlefield operations, and autonomous driving typically require a focus on moving objects. In such scenarios, acquiring information about the motion state of objects in relation to the ground poses a challenge that current methods find difficult to tackle. Furthermore, as many scenarios utilize single RGB cameras rather than stereo or RGB-D cameras, discerning object motion becomes even more challenging. Addressing this issue, this paper introduces a straightforward yet practical approach for determining object motion, adaptable across various environments.
@@
Traditional non-deep learning methods adeptly track moving targets, harnessing the physical and mathematical facets of images and relying on comprehensive mathematical models. While they might exhibit reduced accuracy in select cases, their efficacy extends across almost any environment. In contrast, deep learning-driven MOT frameworks pivot on data-driven principles. Though their performance is notable, they may falter when confronted with limited or inadequate data, leading to inaccuracies in tracking object motion estimation.
To transcend the limitations inherent in both deep learning algorithms and traditional methods, we present a tracking approach that melds conventional mathematical modeling with data-driven deep learning methodologies. In essence, this paper unfolds with three core contributions:
  • 1)
    introducing the Motion State Validation F1 Score (MVF1) as a metric to assess the model’s proficiency in determining the relative motion status of objects in relation to the ground. The experimental validation of MVF1 is conducted, offering substantiating evidence for its soundness;
  • 2)
    presenting a plug-and-play method for discerning the motion attributes of tracked objects;
  • 3)
    proposing a model-data-driven tracking framework(MoD2T) that harmonizes traditional methods with MOT tracking frameworks, enabling the delineation of object motion states.
@@ II-A Motion Objects Detection Methods
Traditional MOT methods primarily focus on identifying moving targets, and the subsequent tracking process is similar to the SORT[4]. In the early stages, image change detection methods were used for this purpose.
The development of image change detection methods can be traced back several decades. The earliest methods were based on pixel value comparison to detect changes [17, 1, 18]. These methods were simple and intuitive but sensitive to complex scenes and noise.
Subsequently, methods based on pixel-level differences received extensive research attention. These methods use pixel-level statistics or thresholds to detect changes, such as pixel ratio [19] and statistical parameters [20]. While these methods are robust to noise, they still struggle with complex textures and lighting variations [21].
Before the surge of deep learning, various Gaussian background model were the state-of-the-art in this field. Specifically, for a background image, the grayscale values of specific pixel points in the temporal domain follow a Gaussian distribution. In other words, for a background image $B$, the grayscale value $x$ of each point $(u,v)$ satisfies $I(x)\sim N\left(\mu,\sigma^{2}\right)$ as shown in (1):
@@
If the condition is met, the point is considered a background point; otherwise, it is considered a foreground point. Here, $T$ is a constant threshold representing the gray scale value at a given point in image $IG$ and $IB$ is similar in nature.
@@
In the field of image processing, where deep learning methods dominate, motion detection methods primarily adopt sparse feature tracking[22, 23, 24] or dense optical flow [25, 26, 27]. The former is widely used in Simultaneous Localization and Mapping(SLAM), while the latter is commonly applied for overall motion estimation in images. Although each method has been proven effective for its respective applications, neither can fully model the motion in videos: optical flow cannot capture long-term motion trajectories, and sparse tracking cannot model the motion of all pixels. Recently, someone proposed a method combining the Optical flow with some other modes [28], and achieved good results.
However, these methods still have certain limitations. Sparse feature tracking involves random selection of feature points, which may result in tracking objects with too few or no features, or unnecessary computational overhead if too many features are selected. The limitations of optical flow methods based on CNN[29, 30] or generative adversarial networks (GAN)[31, 32] include current challenges regarding their accuracy and universality. Traditional optical flow methods, on the other hand, exhibit high tracking accuracy in current MOT tracking frameworks.
Of course, in recent years, several models based on deep learning for detecting changes in images have been proposed [33, 34, 35]. These methods have been widely used in fields such as remote sensing and change detection. However, due to the fact that their working principle relies on static image comparison, these methods are no longer suitable for motion object detection.
II-B MOT Methods And Fusion Methods
The developmental trajectory of MOT methodologies has been expounded upon in the preceding text and will not be extensively reiterated here.
Currently, the available MOT approaches capable of distinguishing object motion exhibit limited proliferation. A significant portion of these approaches do not solely rely on RGB image information, but rather necessitate supplementary sensing modalities such as RGB-D cameras[36] or various alternative sensors like LIDAR[22].
The conventional MOT approach employed within this paoer shall be elucidated in the appendix. As for the deep learning-driven MOT method, this paper intends to embrace the tracking-by-detection paradigm.
For the object detection stage, YOLOv8[37] will be used. YOLOv8 is a fast and accurate object detection method that divides the image into grids and predicts bounding boxes and their corresponding classes for object detection.
For the subsequent tracking stage, StrongSORT[38] will be employed. StrongSORT is a deep learning-based object tracking method that combines appearance features and motion information to achieve accurate multiple object tracking.
The tracking framework, YOLOv8+StrongSORT, combines the YOLOv8 object detector with the StrongSORT tracker, resulting in powerful tracking performance. YOLOv8 can rapidly and accurately detect targets in video frames and provide bounding box locations and class information for each detection. StrongSORT utilizes these detection results and combines them with the appearance features and motion information of the targets. It performs data association and trajectory prediction for target tracking and re-identification.
Additionally, the paper has also experimented with other tracking methods, such as YOLOv8+ DeepOCSORT [6], to explore multiple tracking approaches. Our method can be integrated with different MOT methods, and this article only selected two.
@@ II-B MOT Methods And Fusion Methods
Figure 1: Flow chart of motion state judgment. The background feature extraction process is shown in Figure 2. And, in this paper, The frame rate difference $n$ for calculating distance and background similarity is set to 3. The specific meaning of the formula and the specific meanings of each parameter can be found in section III. The main purpose of this process is to calculate two indicators, one is the distance of the object’s movement in the image, and the other is the similarity of the object’s background in the image. Finally, these two indicators are combined to determine whether the object is moving relative to the ground.
The fusion of tracking methods involves combining multiple individual tracking methods or models to improve tracking performance or address various challenges. This approach can leverage the strengths of different methods, mitigate the limitations of individual methods, and provide more robust and accurate tracking results.
One common method of tracking method fusion is decision-level fusion [39, 40, 41]. This method determines the final tracking result by weighting or voting the outputs of multiple trackers. The weights of each tracker can be dynamically adjusted based on its tracking performance in previous frames, giving greater weight to trackers with better performance.
Another common fusion method is feature-level fusion [42, 43, 44, 45, 46]. In this approach, multiple trackers extract feature representations of the target and fuse these features. Fusion can be as simple as averaging or weighted averaging, or it can involve more complex model fusion methods such as convolutional neural networks or support vector machines.
Additionally, there are information fusion-based methods [47, 48] that involve fusing motion information, contextual information, or external prior knowledge. These methods improve tracking performance by combining information from multiple sources. For example, the results of multiple trackers can be fused by incorporating the target’s motion model, scene context, or appearance model.
Overall, the fusion of tracking methods offers a way to enhance tracking performance by integrating multiple methods or models, leveraging their strengths, and taking advantage of different sources of information.
III Motion State Judgment
III-A Fusion Method
@@
III-B Quantifying Object Motion
MOT methods are employed to discern the motion trajectories of objects within video sequences. In scenarios involving camera motion, the application of the enhanced correlation coefficient maximization(ECC) [49] model enables robust camera motion correction, allowing us to reliably ascertain the relative inter-frame displacement of objects. This computed displacement serves as a foundational metric for characterizing object motion.
We employ the Mahalanobis distance to calculate the motion distance $D$, as shown in (1):
@@ III-B Quantifying Object Motion
Here, $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$ represent the center coordinates of the two bounding boxes, and $(x_{1},y_{1},w_{1},h_{1})$ and $(x_{2},y_{2},w_{2},h_{2})$ denote the coordinates and dimensions of the two bounding boxes.
The rationale behind selecting Mahalanobis distance lies in its ability to account for the covariance structure within the data. This becomes particularly crucial when dealing with multidimensional data, as it considers both the variances and covariances among the different dimensions. Notably, we conducted experiments with alternative distance metrics, and Mahalanobis distance demonstrated superior performance in capturing the underlying data structure, especially in handling the intricacies of motion patterns represented by the bounding boxes. This choice significantly contributes to the accuracy and reliability of our motion distance calculations, ensuring robustness in the face of varying data distributions and dimensions.
Figure 3: The comprehensive architecture of MoD2T is presented, allowing for flexibility in the integration of various deep learning-based methods. Specifically, this study employs YOLOv8+StrongSORT and YOLOv8+DeepOCSORT as instances of such methods. Figure.7 illustrates the traditional method, while Figure.1 focuses on the motion state judgment aspect. Within the algorithmic effectiveness assessment module depicted in the figure, the region enclosed by purple dashed lines signifies multiple potential outputs of this module. The notation $p_{t}$ within the diagram represents the pixels that traditional algorithms may identify as moving objects.
III-C Background Feature Extraction
We performed feature extraction on the background of the tracked object. The specific approach involved segmenting the surrounding background and then combining it. We utilized a pre-trained OSNet [50] to extract appearance features from this background. The procedure is illustrated in the Figure.2.
Empirical findings have substantiated the enhancement in accuracy resulting from the incorporation of this procedural step. The discernible rise in MVF1 scores, as illustrated in Tables IV, V, and VI, under the condition $\lambda=best$, underscores its efficacy in contrast to cases where $\lambda$ is set to 0. The subsequent section will provide a comprehensive explication of the variable $\lambda$, elucidating its impact on MVF1 scores. The optimal value for $\lambda$, denoted as $best$, is determined through systematic parameter tuning, with specific values tailored for each dataset. For brevity, explicit enumeration of these dataset-specific $best$ values is omitted in this context.
III-D Motion State Judgment